OpenStudy (anonymous):

Which of the following represents eighth root of x cubed in exponential form?

OpenStudy (anonymous):

Here's a start: \[\sqrt[8]{x^{3}} = (x^{3})^{\frac{ 1 }{ 8 }}\]

OpenStudy (anonymous):

so would this be the answer x to the 8 thirds power

OpenStudy (anonymous):

??

OpenStudy (anonymous):

why eight thirds? when you have exponents like this, you multiply the exponents. \[(x^{a})^{b} = x^{ab}\]

OpenStudy (anonymous):

so it would be ?

OpenStudy (anonymous):

so if a was 3 and b was (1/8), then x^ab would be x^(3/8) \[(x^{3})^{\frac{ 1 }{ 8 }} = x^{\frac{ 3 }{ 8 }}\]